DVICOPY(1) Web2C7.32x DVICOPY(1) NAME dvicopy−producemodifiedcopyofDVIfile SYNOPSIS dvicopy[options][infile[outfile]] DESCRIPTION Thismanualpageisnotmeanttobeexhaustive.Thecompletedocumentationforthisversionof canbefoundintheinfofileormanualWeb2C:ATeXimplementation. dvicopyreadsaDVIfile,expandsanyreferencestovirtualfontstobasefonts,andwritesthe resultingDVIfile.ThusyoucanusevirtualfontsevenifyourDVIprocessordoesnotsupport them,bypassingthedocumentsthroughdvicopyfirst. SEEALSO dviselect(1) 24January1999 TeX 1